OpenStudy (anonymous):

As a salesperson at Roaring Waves Beach Supplies, Carissa receives a monthly base pay plus commission on all that she sells. If she sells $500 worth of merchandise in one month, she is paid $390. If she sells $1,000 of merchandise in one month, she is paid $480. Find Carissa's salary when she sells $2,400 worth of merchandise. @Mertsj

OpenStudy (mertsj):

Well...we don't know her base pay and we don't know the commission rate so we need two variables

OpenStudy (mertsj):

If her base pay is b and her commission percent is r then her pay would be b+br for each week.

OpenStudy (mertsj):

If her base pay is b and her commission percent is r then her pay would be b+r(500) for the first month.

OpenStudy (mertsj):

And for the next month it would be b+r(1000) for the next month.

OpenStudy (mertsj):

So now we have two equations: the first month: b+500r=390 the second month: b+1000r=480

OpenStudy (anonymous):

Ok.

OpenStudy (mertsj):

So subtract the first equation from the second equation and we have: 500r = 90 and r = .18

OpenStudy (mertsj):

So now substitute back into one of the equations and find b b+1000(.18)=500 b+180=500 b=320

OpenStudy (mertsj):

So if she sells $2400 worth of stuff she will get 320 +.18(2400)

OpenStudy (anonymous):

So do I solve that last part? 320 +.18(2400)?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

752 is what I got.

OpenStudy (mertsj):

Me too.

OpenStudy (anonymous):

Okay :) so that's the final answer?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

Was wrong :l.. Don't understand lol...

OpenStudy (mertsj):

Let me double check.

OpenStudy (anonymous):

Ok

OpenStudy (mertsj):

Try 732

OpenStudy (anonymous):

Correct :)! thanks!!!

OpenStudy (mertsj):

yw